If you find that [CoCl42-(aq)] = 4.431 M, [Co(H2O)62+(aq)] = 0.172 M, and [Cl-(aq)] = 0.757 M at 7 ∘C, what is value of ΔG at 7 ∘C, in kJ/mol?
Answer #1

The expression for the equilibrium constant (K) can be written as follows.

K = [CoCl42-(aq)]/[Co(H2O)62+(aq)][Cl-(aq)]4

i.e. K = 4.431/(0.172*0.7574) = 78.45

Now, \DeltaG = -RTln(K)

i.e. \DeltaG = -8.314*10-3 kJ/mol.K * (7+273) K * ln(78.45)

Therefore, \DeltaG = -10.155 kJ/mol